### v2

# Magneto-optical Kerr effect in an A-type antiferromagnet

Abstract
Magneto-optic Kerr effect (MOKE) is a powerful probe of broken time-reversal symmetry ($\mathcal{T}$), typically used to study ferromagnets. While MOKE has been observed in some antiferromagnets (AFMs) with vanishing magnetization, it is often associated with structures whose symmetry is lower than basic collinear, bipartite order. In contrast, theory predicts a mechanism for MOKE intrinsic to all AFMs of A-type, i.e. layered AFMs in which ferromagnetic layers are antiferromagnetically aligned. Here we report the first experimental confirmation of this mechanism in a bulk AFM. We achieve this by measuring the imaginary component of MOKE as a function of photon energy in MnBi2Te4, an A-type AFM where $\mathcal{T}$ is preserved in combination with a translation. By comparing the experimental results with model calculations, we demonstrate that observable MOKE should be expected in all collinear A-type AFMs with out-of-plane spin order, thus enabling optical detection of AFM domains and expanding the scope of MOKE to few-layer AFMs.
The detection of broken time-reversal symmetry ($\mathcal{T}$) in quantum materials is of fundamental interest and practical relevance. Recent advances in the synthesis and exfoliation of van der Waals magnets have highlighted the critical role of optical techniques in detecting $\mathcal{T}$-breaking in samples just a few atomic layers thick [17, 25, 31, 14, 16, 26, 51, 42, 37], whose small mass poses a challenge for bulk probes of magnetism. The magneto-optic Kerr effect (MOKE), which is the difference in reflectivity for left and right circularly polarized (LCP and RCP) light, is routinely used for detection of ferromagnetic (FM) order. Which classes of antiferromagnetic (AFM) order yield a measurable MOKE signal, and the underlying mechanism in each case, are questions of long-standing fascination. Here we show that MOKE is a more powerful and general probe of AFM order than previously thought, with implications for research and applications of both bulk and few-layer AFMs.
Time reversal flips the direction of spin, turning a ferromagnet into its time-reversed counterpart (Fig. 1a). The two configurations with opposite magnetization $M$, called domains, enable the storage, processing, and retrieval of information in the form of a classical 0 or 1. Antiferromagnetic (AFM) order is also characterized by degenerate ground states related by $\mathcal{T}$, but $M=0$ in each domain. In the simplest example neighboring spins are antiparallel (Fig. 1b). AFMs play an integral role in critical technologies [4, 33], and interest in their properties has continued to grow [19, 8, 41, 43]. Zero magnetization can be a feature, as the absence of the long-range magnetic dipolar interaction leads to enhanced scalability.
Figure 1: Schematics of representative magnetic structures. (a) Ferromagnet, where $\mathcal{T}$ reverses the sign of the order parameter; (b) A-type antiferromagnet, where the sign reversal by $\mathcal{T}$ can be undone by a translation by half a magnetic unit cell ($S_{1/2}$); (c) Thin flakes of an A-type antiferromagnet: flakes of odd layer number NN have a net magnetic moment, while those with even NN do not. (d) Two layers of the MnBi2Te4 crystal and magnetic structure, demonstrating the A-type AFM phase. Opposite spins are related by $S_{1/2}$.
A non-zero $M$ induces a difference in the index of refraction of LCP and RCP light, yielding a mechanism for MOKE that is essentially the same in all FMs. MOKE is therefore a reliable technique for distinguishing FM domains with high-spatial resolution. In contrast, domains have been detected through MOKE only in a few AFMs [29, 22, 46], and the mechanism that underlies this signal has been attributed to specific properties of individual compounds.
Here we focus on a class of AFMs in which the time-reversed states are related by translation by half the magnetic unit cell, $S_{1/2}$, illustrated in bulk and thin flake forms in Figs. 1b and c, respectively. We choose this class because the product $\mathcal{T}S_{1/2}$ ensures that the index of refraction of LCP and RCP light is the same as they propagate through an unbounded medium, a condition that at first sight seems to stipulate the absence of MOKE, since it is the difference of the two indices that causes MOKE in FMs. However, this conclusion is incorrect: MOKE by its nature involves a bounding surface that breaks $\mathcal{T}S_{1/2}$, unless $S_{1/2}$ is parallel to the surface. Therefore, $\mathcal{T}S_{1/2}$ imposes no constraints on MOKE at surfaces that break translational invariance described by $S_{1/2}$.
We investigate MnBi2Te4, a layered topological AFM with a Néel temperature of 25 K [35], as an example of a $\mathcal{T}S_{1/2}$ - symmetric AFM. Spins in individual Mn layers in MnBi2Te4 are parallel to each other, and antiparallel to the spins in neighboring layers (Fig. 1d). This order, called A-type, persists to the surface of MnBi2Te4 [40]. MOKE was recently reported in the wavelength range of $500-1000\,\mathrm{nm}$ in thin flakes of MnBi2Te4, in samples with an odd number of layers ($N$), which have a net magnetization, and in samples with an even $N$, which do not [36, 48, 39, 5]. While symmetry permits MOKE for all $N$, it does not identify the underlying mechanism.
Axion electrodynamics associated with topology was proposed as the mechanism for MOKE in MnBi2Te4 [39, 1]. The discontinuity in the topological $\mathcal{Z}_{2}$ invariant at the sample/vacuum interface manifests as a quantized surface Hall conductance, which is indistinguishable from a quantized trace of the magnetoelectric (ME) tensor in the static limit [13]. The surface conductance gives rise to a traceful, non-quantized ME tensor above zero frequency [30, 1]. Optical phenomena associated with a traceful ME tensor (collectively known as axion electrodynamics [44]), provide a mechanism for MOKE [23, 34, 1], as demonstrated in studies of the non-topological ME Cr2O3 [29].
A distinct mechanism for MOKE in $\mathcal{T}S_{1/2}$ symmetric AFMs was proposed by Dzyaloshinskii and Papamichail [12]. They treated the layered AFM as a stack of FMs with alternating spin direction. The dielectric tensor of each layer is assumed to be the same as in an unbounded medium with the same FM order, and MOKE arises due to the variation in dielectric properties as the wave propagates. In the ‘alternating FM’ framework, the optical response arises from the bulk magnetic structure, and the surface conductance does not play a special role, unlike in the axion electrodynamics mechanism, where the surface contribution is essential. To the best of our knowledge, MOKE arising from the alternating FM effect has not been experimentally recognized prior to this work.
Below we report measurements of MOKE in ‘infinite’ ($\sim 136\,\mathrm{nm}$ thick, Fig. ED1) layer MnBi2Te4 as a function of magnetic field, $H$, and wavelength of the optical probe, $\lambda$. We find that non-zero MOKE appears in the $M=0$ state, and that its spectrum is quantitatively captured by the alternating FM model [12], without invoking axion electrodynamics. The alternating FM mechanism is allowed in all collinear AFMs whose spins are oriented perpendicular to the layers, regardless of the number of layers, band topology or symmetry, opening the door for MOKE studies of a broader class of AFMs than was previously considered feasible.
The experiment: Magnetic field, position and wavelength dependence
Figure 2: Magneto-optical measurements of MnBi2Te4. (a) Reflection circular dichroism (RCD) measured at $525\,\mathrm{nm}$ as a function of magnetic field, showing a jump at the spin-flop transition ($H_{SF}\approx 3.4\,\mathrm{T}$) and a hysteresis loop with a coercive field of $H_{C}\approx 1\,\mathrm{T}$. A small setup-induced background contribution has been subtracted (Fig. ED5). RCD is compared with bulk magnetization measurements taken at $15\,\mathrm{K}$ with the field applied along the crystallographic $c$-axis (black line). The inset illustrates the bulk spin-flop transition. (b) Spatially resolved RCD measurements at $H=0$, revealing AFM domain structures that change between cooldowns.
The complex MOKE angle is defined by the relation, $\Theta\equiv-i\delta r/r$, where the reflectivities of incident LCP and RCP light are given by $r=r_{0}\pm\delta r$ (the sign and polarization conventions are defined in the Methods section). Throughout this paper we show measurements of the imaginary part of $\Theta$, which is known as reflection circular dichroism (RCD). We measure RCD using a photoelastic modulator to vary the helicity of the incident light at $50\,\mathrm{kHz}$, using a setup illustrated in Fig. ED2. The change in reflectivity synchronous with the helicity modulation is proportional to RCD (see Methods for details of the experiment).
In Fig. 2a we compare the magnetic field dependence of RCD to that of bulk magnetization at $15\,\mathrm{K}$, measured on a sample from the same batch. RCD reveals two discontinuous features: a jump at $H_{SF}\approx 3.4\,\mathrm{T}$ and a hysteresis loop with a coercive field of $H_{C}\approx 1\,\mathrm{T}$. The jump corresponds to a spin-flop (SF) transition, illustrated in the inset, which is also seen in magnetization. The corresponding magnetization increase is $\Delta M_{SF}\approx 1.2\left(\mu_{B}/\text{Mn}\right)\approx 0.4M_{s}$, where $M_{s}=3\left(\mu_{B}/\text{Mn}\right)$ is the ordered moment at $15\,\mathrm{K}$ [47]. The hysteresis loop has no counterpart in bulk magnetization measurements, which find $M=0$ for $H<H_{SF}$. Therefore, the two non-zero values of RCD at $H=0$ correspond to time-reversed states, each with $M=0$. Further evidence for this interpretation is provided by spatially resolved measurements (Fig. 2b) that reveal spontaneous formation of domains with positive and negative RCD values in zero field, with distinct structures in different $H=0$ cooldowns.
The existence of two discontinuities in RCD vs. $H$ provides an unusual opportunity to compare the MOKE spectrum in a phase with $M\neq 0$, that is the spin-flop (SF) phase, and a $\mathcal{T}S_{1/2}$ -symmetric AFM ($M=0$) phase in the same material. We measured the magnetic field dependence of RCD in the photon energy range from $1.34\,\mathrm{eV}$ ($925\,\mathrm{nm}$) to $3.22\,\mathrm{eV}$ ($385\,\mathrm{nm}$) (Fig. ED6), and extracted RCD spectra in the two phases. In Fig. 3a we compare the RCD spectrum corresponding to the SF transition with the RCD spectrum in the AFM phase. The AFM spectrum is half of the difference between RCD at $H=0^{+}$ and $H=0^{-}$, where $H=0^{+}$ and $H=0^{-}$ denote zero field as approached from positive and negative fields. The SF spectrum, corresponding to the jump in magnetization across the transition, is given by $\Theta_{SF}=\Theta(3.4\,\mathrm{T})-\left[\Theta(3.2\,\mathrm{T})-\Theta(0^{+})\right]$, because $\Theta(3.2\,\mathrm{T})$ is a sum of the antiferromagnetic and magnetization contributions, while in the spin flop phase, at $3.4\,\mathrm{T}$, there is no antiferromagnetic contribution.
The two spectra are strikingly different, ruling out a small uncompensated bulk ferromagnetic moment as a trivial explanation for RCD at $H=0$. Further, the fact that the two spectra are rich in features creates an opportunity to constrain theories: a suitable theoretical model should simultaneously reproduce both of them.
Figure 3: Spectroscopic signature of MOKE in antiferromagnetic MnBi2Te4. (a) Measured RCD spectrum at the spin-flop transition (purple) and at zero field (red), showing distinct spectral shapes. For each photon energy, we measured the magnetic field dependence of RCD (Fig. ED6). The AFM spectrum is half of the difference between RCD on the downward ($H=0^{+}$) and upward ($H=0^{-}$) field sweep, averaged across the hysteresis loop, between $\pm 0.9\,\mathrm{T}$. The SF spectrum is given by $\Theta_{SF}=\Theta(3.4\,\mathrm{T})-\left[\Theta(3.2\,\mathrm{T})-\Theta(0^{+})\right]$. The error bars for the $H=0^{+}$ measurement represent the standard deviation across this field range. The larger error bars for the SF spectrum reflect the fact fewer measurements were used to extract the corresponding RCD value (at $H=3.4\,\mathrm{T}$ and $H=3.2\,\mathrm{T}$). Laser power fluctuations do not contribute to these error bars, since the RCD is normalized by the simultaneously measured reflectivity (the same detector voltage is demodulated at two frequencies). The only relevant systematic uncertainty is from the PEM calibration, which for our Hinds PEM-100 is $\leq 5\%$, corresponding to <0.1<0.1 mrad on the RCD scale.
(b) RCD spectra computed from the Lorentz model (Eq. 1, with parameters from Table ED1) for a ferromagnet (purple, schematic in panel c) and two antiferromagnetic domains (red and blue, schematics in panel d). The ferromagnetic RCD is scaled by a factor of $0.4$ from that calculated for a saturated moment, to be directly comparable to the measured RCD jump at the spin-flop transition.
Lorentz model: “infinite” and “finite” samples
The RCD spectrum measured in the SF state is consistent with resonant MOKE in a FM. Both the magnitude and zero crossing can be reproduced by a Lorentz oscillator whose resonant photon energy differs for LC and RC polarized light. The corresponding dielectric function is:
$\varepsilon_{\pm}(\omega)=\varepsilon_{\infty}+\frac{f}{(\omega_{0}\pm\delta\omega)^{2}-\omega^{2}-i\gamma\omega},$ (1)
where $f$, $\omega_{0}$, and $\gamma$ are oscillator strength, resonant frequency, and damping. These parameters, together with a background contribution, $\varepsilon_{\infty}$, are chosen to reproduce the measured optical conductivity in the spectral range of interest, reported in Ref. [27]. The good agreement between the measured quantities and those calculated from the Lorentz model is shown in Fig. ED3. Finally, $\delta\omega$, the difference in resonance frequency for LC and RC polarized light, is chosen to fit the magnitude of the measured RCD. The purple curve in Fig. 3b shows that the single oscillator model captures the main features of the observed ferromagnetic RCD spectrum (Fig. 3a).
Our quantitative description of RCD in the ferromagnetic phase of MnBi2Te4 leads us to consider the alternating FM model for the RCD in its AFM state: a stack of FM layers in which the sign of $\delta\omega$ alternates in neighboring layers. We use the transfer matrix formalism [49], following the approach in Ref. [21], to calculate the reflection coefficient from the stack. The number of layers is 100, and the thickness of each layer, $d=\mathrm{1.3}{nm}$, corresponds to the separation between spins in MnBi2Te4. The simulated RCD spectra for the two $\mathcal{T}$-reversed AFM domains (Fig. 3d), are shown in red and blue in Fig. 3b. Remarkably, this minimal model captures the spectral shape and the magnitude of RCD in the AFM phase of MnBi2Te4, without introducing free parameters beyond those that describe the FM phase. In the SI we extend this analysis by using a dielectric function computed from a density functional theory (DFT) calculation to construct the alternating FM model, and we find that this approach also yields the correct magnitude of RCD in the AFM phase (Fig. ED4). To the best of our knowledge, this is the first experimental evidence for an RCD arising from an alternating FM model in a bulk AFM.
Figure 4: Dependence of circular dichroism on the number of layers (a, b) Calculated reflection (RCD) and transmission (TCD) circular dichroism as a function of layer number for even and odd layer stacks, using the same model and parameters as in Fig. 3. (c, d) RCD and TCD spectra for 8-layer (8L) and 9-layer (9L) AFM stacks.
Next, we show that the key features of RCD and transmission circular dichroism (TCD) that are seen in few layer crystals in Ref. [39] can be understood within the same model. We performed a series of revealing transfer matrix calculations, varying the number of layers while keeping the thickness of individual layers fixed. In Figs. 4(a,b) we plot the calculated RCD and TCD, respectively, as a function of layer number at fixed photon energy. We find that both RCD and TCD depend strongly on $N$ if it is odd, but are independent of $N$ when it is even, consistent with the findings of Ref. [12]. When the slab thickness $Nd$ far exceeds the optical penetration depth, RCD in even and odd layer samples converge to the same value, as expected. TCD vanishes for all even layer stacks, a phenomenon that can be understood from a symmetry perspective: even layer stacks are symmetric under a product of inversion and time-reversal symmetry, which prohibits circular dichroism in transmission [11, 7, 3, 39].
Further, the dichroism spectra for even and odd layer stacks are strikingly different (Fig. 4c, d). For $N=8$, the RCD spectrum is identical to the spectrum of the bulk AFM crystal (Fig. 3b), while spectrum for $N=9$ resembles that of the bulk SF phase, as $M\neq 0$ for odd layer stacks. The maximum RCD is about three times larger for $N=9$ than for $N=8$. However, these calculations assume that samples are suspended in vacuum, which is not a realistic experimental scenario. The transfer matrix approach allows us to include the substrate in the calculation of reflectivity [21], and we find that doing so preferentially suppresses dichroism in odd-layer samples. In Fig. ED7 we show that the maximum magnitude of RCD for samples on sapphire and diamond substrates, used in Ref. [39], is approximately equal in magnitude for even and odd layer stacks, while they maintain the characteristic spectral shapes shown in Fig. 4c. We therefore find that the alternating FM model with realistic parameters captures the main experimental results on $N$-even and $N$-odd flakes of MnBi2Te4 reported in [39], as well as on ‘infinite’ layer flakes reported here.
Bilayer: the exact solution and the physical picture
While the numerical results shown in the above section capture the experimental results, the underlying physics is somewhat mysterious, and poses the following questions: why does MOKE arise in $\mathcal{T}S_{1/2}$-symmetric AFMs within the alternating FM model, and what influences its magnitude? Why is its value independent of the (even) number of layers? Why does TCD vanish? To address these questions, we first consider the analytical solution for MOKE in a bilayer, and then offer an intuitive physical picture, shedding light on the experimental, numerical and analytical findings.
We consider a single bilayer of oppositely oriented FMs. Each layer is characterized by the index of refraction $n+\delta n$, where the sign of $\delta n$ is opposite for the two layers (Fig. 5a). First, we utilize the transfer matrix approach to reproduce the expressions for the transmission and reflection coefficients of such a bilayer, derived in Ref. [12]. MOKE in the physically relevant limit ($k_{0}d\ll 1$ and $\delta n\ll n$) is given by
$\Theta_{AFM}=-k_{0}nd\frac{2\delta n}{n^{2}-1}=\left(-ik_{0}nd\right)\Theta_{FM},$ (2)
where $\Theta_{FM}$ is the Kerr angle of the corresponding bulk ferromagnet, $k_{0}=\omega/c$ the wavevector in vacuum, and $n$ the average index of refraction of the two layers. The expression for arbitrary layer thickness and $\delta n$ is given in the Supplementary Information.
Eq. 2 showcases the common microscopic origin of the MOKE response of an AFM bilayer and that of the corresponding ferromagnet, but it also reveals a surprising aspect of the link between them. It is natural to assume that a $\mathcal{T}S_{1/2}$ - symmetric type-A AFM can exhibit nonzero MOKE only if strong absorption (governed by the imaginary part of $n$) leads to disproportionate sampling of the top layer. However, Eq. 2 shows that this is not correct: the magnitude of $\Theta_{AFM}$ depends on the absolute value of the index, $|n|$, and is nonzero even in the absence of dissipation. The phase of $n$ appears in the phase shift between $\Theta_{AFM}$ and $\Theta_{FM}$, which manifests as the difference in the spectrum of RCD between AFM and FM structures, as seen for example in Fig. 3a.
Figure 5: Illustration of the model for MOKE in an antiferromagnetic bilayer. (a) Schematic representation of a bilayer, consisting of two layers of indices of refraction $n_{1,2}=n\pm\delta n$ when probed with circularly polarized light. $\delta n$ is induced by the opposite magnetization in the two layers, $\pm M_{z}$. (b,c) Simplified representation of a bilayer, where the layers are replaced by two infinitesimally thin sheets of conductances $G_{1,2}=G_{0}\pm\delta G$, separated by a dielectric medium of thickness $d$ and index of refraction $n$. The propagation-induced differences between fields emitted from the two sheets cancel in (b) transmission, but not in (c) reflection, leading to vanishing TCD and non-zero RCD.
To gain further insight into the physical mechanism behind the alternating FM scenario, we analyze a toy model (Fig. 5(b,c)) in which each of the two layers is replaced by an infinitesimally thin ferromagnetic sheet of optical conductance $G_{1,2}$, where $1,2$ are layer indices. The sheets are separated by a non-magnetic medium of thickness $d$ and index of refraction $n$. For illustrative purposes we analyze a single reflection from each sheet.
We consider a circularly polarized plane wave propagating in the positive $z$ direction with amplitude $E_{0}$. The electric field generates sheet currents,
$K_{1,2}=G_{1,2}E_{0}e^{\mp ik_{0}nd/2},$ (3)
where we have taken $z=0$ as the midpoint between the sheets for convenience. The currents $K_{1,2}$ radiate, emitting electric fields,
$E_{1,2}(z)=\frac{\mu_{0}c}{2}K_{1,2}e^{ik_{0}\left|z\pm d/2\right|},$ (4)
where $c$ is the speed of light and $\mu_{0}$ the vacuum permeability.
The total transmitted electric field, given by the sum of $E_{1}(z)$ and $E_{2}(z)$ for $z>d/2$, is,
$E_{T}(z)=\mu_{0}cE_{0}\left(\frac{G_{1}+G_{2}}{2}\right)e^{ik_{0}z}=\mu_{0}cE_{0}G_{0}e^{ik_{0}z}.$ (5)
As $E_{T}(z)$ depends only on the average conductance $G_{0}$, the TCD is zero. The toy model shows that this is a consequence of cancellation of propagation factors. The incident wave propagates an additional distance $d$ in the medium beyond layer 1 to induce currents layer 2. However, the wave emitted by layer 1 must propagate through the medium to reach a given point $z>d/2$. The total propagation factor, $e^{iknz}$, for waves transmitted through the bilayer is the same, and independent of $d$ (Fig. 5b). The sum of fields emitted from the two sheets therefore depends only on the sum of the currents, $G_{1}+G_{2}$.
In contrast, the propagation factors for fields radiated in the z-z direction add, rather than cancel: the incident EM wave travels further to excite current in layer 2 and its radiated field propagates further before reaching an observer at z<d/2z<-d/2 (Fig. 5c). The total reflected electric field is given by,
$E_{R}(z)=$ $\frac{\mu_{0}cE_{0}}{2}e^{-ik_{0}z}\left(G_{1}e^{-ik_{0}nd}+G_{2}e^{ik_{0}nd}\right)$ (6)
$\approx\mu_{0}cE_{0}e^{-ik_{0}z}\left(G_{0}-\delta Gik_{0}nd\right),$
where we used $|k_{0}nd|\ll 1$ in the expansion of the exponential. The term proportional to $\delta G$ gives rise to $\Theta_{AFM}$.
The toy model for a bilayer identifies the origin of nonzero $\Theta_{AFM}$ as the change in the electric field as it propagates, regardless of whether it is dominated by the real or imaginary component of $n$. We see that absorption does not play a special role in generating MOKE. Further, RCD does not depend on the layer number for even layers (Fig. 4a) because of the exponential nature of propagation: the ratio of fields in two neighboring layers is $\exp{(ik_{0}nd)}$, independent of the number of layers. Thus the effects that yield RCD in two atomic layers and in a semi-infinite AFM crystal are the same. These qualitative features found in the toy model are retained when multiple reflections are included through the transfer matrix approach (Eq. 2, and Fig. 4).
Conclusion
In this paper we identify alternating FM layers as the origin of MOKE in an A-type antiferromagnet with $\mathcal{T}S_{1/2}$ symmetry, therefore providing first experimental confirmation of the mechanism proposed by Dzyaloshinskii and Papamichail [12]. Within this theory, $\Theta_{AFM}$ is entirely determined by the MOKE response of the ferromagnetic layers that are the building blocks of the AFM, allowing us to test it directly by tuning MnBi2Te4 across the spin-flop transition, thus accessing $\Theta_{AFM}$ and $\Theta_{FM}$ in the same crystal. Based on the quantitative agreement between the measured RCD spectrum in the AFM phase and the theoretical prediction based solely on the FM spectrum, we conclude that MOKE observed at visible and near infrared wavelengths in MnBi2Te4 arises from alternating FM layers, rather than axion electrodynamics associated with band topology. Axion electrodynamics may dominate the MOKE response at terahertz frequencies below the bulk band gap [30, 1, 20, 45], where there are no other optical transitions that could overwhelm their contribution, although results in even-layer flakes suggest that the axion contribution in this frequency range is weak [30, 20].
The effect we observe is not limited to MnBi2Te4, and is generically present in A-type antiferromagnets with spins out of plane. We believe that it has not been recognized experimentally in part because the signal vanishes when averaged over atomic steps, naturally present in as-grown bulk crystals. However, optical experiments in van der Waals materials regularly probe surfaces that are atomically flat across the length scale of a laser beam. We note that some previous sightings of MOKE in van der Waals AFMs with an even number of layers were conditionally attributed to extrinsic magnetization [24, 50, 48], given the widespread conviction that MOKE intrinsic to A-type AFMs was forbidden by symmetry. While this explanation is certainly plausible, our findings show that intrinsic MOKE in A-type AFMs is always allowed, and other mechanisms should be considered when $\Theta_{AFM}$ cannot account for the magnitude, or the spectrum, of the observed signal.
Correctly estimating $\Theta_{AFM}$ is therefore essential for interpretation of experimental results. Our work suggests a strategy to do so, both in bulk and thin flakes. In bulk, it is sufficient to combine measured $\Theta_{FM}$ and optical conductivity with the analytical expression for $\Theta_{AFM}$ (Eq. 2). For thin flakes, it is important to consider the role of the substrate, which may suppress or enhance observed signals [21]; we suggest the transfer matrix approach as a convenient way to do so. We note that $|\Theta_{AFM}|$ is not generically small compared to the noise floor of modern experimental setups: it is suppressed with respect to $|\Theta_{FM}|$ only by a factor of $|k_{0}nd|=|2\pi nd/\lambda|$, which is more than an order of magnitude larger than $d/\lambda$ suggested by dimensional analysis.
Our findings show that MOKE is a powerful and general tool for investigating $\mathcal{T}S_{1/2}$ symmetric AFMs. In this work we used it to image time-reversed domains, as shown for in Fig. 2b.
Further, the nonzero $\Theta_{AFM}$ enabled the discovery that a magnetic field can select antiferromagnetic domains (Fig. 2a) in such magnets, challenging the conventional view that antiferromagnets cannot be controlled by external magnetic fields. Uncovering the mechanism behind field control is beyond the scope of this study, but we note that the weakened exchange interaction at the surface is a plausible cause [40, 9]. This implies that details of the surface can affect the coercive field, as confirmed by measurements on two flakes exfoliated from the same batch (Fig. S1). However, the magnitude of the RCD at $H=0$ is a bulk property, independent of local surface conditions. The coercivity and field-switching process will be the subject of future computational and experimental investigations, enabled by the discovery of the AFM RCD signal. In summary, our results address fundamental questions concerning the origin of MOKE in AFMs, provide a strategy for studying and controlling their domains, and elucidate results on exfoliated van der Waals AFMs.
Acknowledgements.
We thank Christine Kuntscher for providing optical conductivity and reflectance data published in Ref. [27], and Nicola Spaldin, Joel Moore and Bevin Huang for useful discussions. V.S. and J.O. received support from the Gordon and Betty Moore Foundation’s EPiQS Initiative through Grant GBMF4537 to J.O. at UC Berkeley. Experimental and theoretical work at LBNL and UC Berkeley was funded by the Quantum Materials (KC2202) program under the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC02-05CH11231. Work at the University of Kansas was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, EPSCoR, and Materials Sciences and Engineering Division under Award No. DE-SC0025319. Parts of device fabrication were performed in the KU Nanofabrication Facility, which is supported by the National Institutes of Health NIGMS P30GM145499. Work at ORNL was supported by the U. S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. For the DFT calculations we used resources provided by the Swedish National Infrastructure for Computing (SNIC) at C3SE. We acknowledge support from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of Office of International Science and Engineering (OISE). This publication is funded in part by a QuantEmX grant from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616 to Sivaloganathan Kumaran.
Supplementary Information for ‘Universal Kerr effect in an A-type antiferromagnet’
The analytical solutions
Bulk ferromagnet
The reflectivity matrix of a ferromagnet in the $\mathbf{\hat{e}}_{\pm}$ basis is:
$r^{FM}_{C}$ $=\begin{pmatrix}r_{0}+\delta r&0\\ 0&r_{0}-\delta r\end{pmatrix}=\begin{pmatrix}\frac{1-n-\delta n}{1+n+\delta n}&0\\ 0&\frac{1-n+\delta n}{1+n-\delta n},\end{pmatrix}$. (S1)
where $n_{1,2}=n\pm\delta n$ are the indices of refraction for $\mathbf{\hat{e}}_{+}$ and $\mathbf{\hat{e}}_{-}$, that is for LCP and RCP incident light, respectively. This leads to:
$\delta r=\frac{1}{2}\left(\frac{1-n-\delta n}{1+n+\delta n}-\frac{1-n+\delta n}{1+n-\delta n}\right)=\frac{-2\delta n}{-\delta n^{2}+n^{2}+2n+1}$ (S2)
and
$r_{0}=\frac{\delta n^{2}-n^{2}+1}{-\delta n^{2}+n^{2}+2n+1},$ (S3)
yielding the complex Kerr angle (Eq. M6):
$\Theta_{FM}=(-i)\left(-\frac{2\delta n}{\delta n^{2}-n^{2}+1}\right)\approx\frac{-2i\delta n}{n^{2}-1}.$ (S4)
Antiferromagnetic bilayer
We use the transfer matrix method to analytically solve for the reflection and transmission of an antiferromagnetic bilayer suspended in vacuum (Fig. 5a). The indices of refraction for $\mathbf{\hat{e}}_{+}$ polarization of the two layers $n_{1,2}=n\pm\delta n$, and $n_{1,2}=n\mp\delta n$ for $\mathbf{\hat{e}}_{-}$. The thickness of the individual layers is $d$, and $k_{0}=\omega/c$, were $\omega$ and $c$ are the frequency and speed of light.
We formulate the calculation in the basis of circular polarization, so that the problem is diagonal. We find that the transmission circular dichroism vanishes identically. In contrast, for reflectivity we find:
$\frac{\delta r}{r_{0}}=\frac{2i\,\delta n\,n\left(\cos\left(2dk_{0}n\right)-\cos\left(2dk_{0}\delta n\right)\right)}{n(-\delta n^{2}+n^{2}-1)\sin\left(2dk_{0}n\right)+\delta n(-\delta n^{2}+n^{2}+1)\sin\left(2dk_{0}\delta n\right)}$ (S5)
The complex MOKE angle can be found using Eq M6. In the physically relevant limit of $dk_{0}\ll 1$, $\delta n\ll n$, this expression reduces to
$\Theta_{AFM}=(-i)\frac{\delta r}{r_{0}}\approx-\frac{2\delta n}{n^{2}-1}(dk_{0}n)=-\frac{2i\delta n}{n^{2}-1}(-idk_{0}n)=\Theta_{FM}(-idk_{0}n)$ (S6)
where $\Theta_{FM}$ stands for the complex Kerr angle of a bulk ferromagnet described by the $n\pm\delta n$ for $\mathbf{\hat{e}}_{+}$.
Photoelastic modulator and the RCD measurement
The time-dependent effect of the photoelastic modulator (PEM) can be captured in the basis of linear polarization with the following Jones matrix:
$J_{PEM}=\left(\begin{array}[]{cc}e^{-\frac{1}{4}i\pi\sin(t\omega_{P})}&0\\ 0&e^{\frac{1}{4}i\pi\sin(t\omega_{P})}\\ \end{array}\right),$ (S7)
where $\omega_{P}$ is the angular frequency of polarization modulation. The modulated light is reflected from the sample, and focused onto a photodiode. The modulation of total reflectivity that is synchronous with the polarization modulation is proportional to the imaginary part of $\Theta$, which can be seen from a Jones matrix calculation, outlined below.
The reflected electric field is given by,
$\begin{pmatrix}E_{x}\\ E_{y}\end{pmatrix}=r_{L}\times J_{PEM}\times\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ 1\end{pmatrix}$, (S8)
where $r_{L}$ is the sample reflectivity expressed in the linear basis. The measured quantity is intensity,
$I=\left|E\right|^{2}\left[\frac{1}{2}\left(1+\operatorname{sgn}\left(\sin{(\omega_{c}t)}\right)\right)\right],$ (S9)
where the square wave term in the square brackets accounts for the modulation by the mechanical chopper at frequency $\omega_{c}$, and $\left|E\right|$ is the absolute value of the reflected electric field. Keeping only the first term in the Fourier transform of the square wave, we have:
$I=\left|E\right|^{2}\left[\frac{1}{2}+\frac{2}{\pi}\sin{(\omega_{c}t)}\right].$ (S10)
To calculate $\left|E\right|^{2}$, we use Eqs. M5, S7 and S8, and find:
$\frac{\left|E\right|^{2}}{\left|r_{0}\right|^{2}}=1+\left|\frac{r_{xy}}{r_{0}}\right|^{2}+2\operatorname{Im}\left(\frac{r_{xy}}{r_{0}}\right)\sin\left(\frac{1}{2}\pi\sin\left(\omega_{P}t\right)\right)\approx 1+4J_{1}\left(\frac{\pi}{2}\right)\operatorname{Im}\left(\Theta\right)\sin\left(\omega_{P}t\right),$ (S11)
where $J_{1}\left(\pi/2\right)$ is the Bessel function of the first kind. In the last step we used the Jacobi-Anger expansion, and omitted the quadratic term in $r_{xy}/r_{0}$. The key point is that this signal contains information on $\operatorname{Im}\left(\Theta\right)$.
The intensity can be expressed as:
$\frac{I}{\left|r_{0}\right|^{2}}=\left(1+4J_{1}\left(\frac{\pi}{2}\right)\operatorname{Im}\left(\Theta\right)\sin\left(\omega_{P}t\right)\right)\left[\frac{1}{2}+\frac{2}{\pi}\sin{(\omega_{c}t)}\right]$ (S12)
$=\frac{1}{2}+\frac{2}{\pi}\sin{(\omega_{c}t)}+2J_{1}\left(\frac{\pi}{2}\right)\operatorname{Im}\left(\Theta\right)\sin\left(\omega_{P}t\right)+\frac{8}{\pi}J_{1}\left(\frac{\pi}{2}\right)\operatorname{Im}\left(\Theta\right)\sin\left(\omega_{P}t\right)\sin{(\omega_{c}t)}.$ (S13)
We simultaneously demodulate the measured intensity at $\omega_{c}$ and $\omega_{P}$ using a lock-in amplifier, yielding currents $I_{c}$ and $I_{P}$, respectively. We find:
$\frac{I_{P}}{I_{c}}=\pi J_{1}\left(\frac{\pi}{2}\right)\operatorname{Im}\left(\Theta\right),$ (S14)
or, equivalently,
$\operatorname{Im}\left(\Theta\right)=\frac{I_{P}}{I_{c}}\frac{1}{\pi J_{1}\left(\pi/2\right)}\approx\frac{I_{P}}{I_{c}}\frac{1}{1.78073}.$ (S15)
Effect of Surface Quality on Coercive Field and Zero-Field RCD
The coercive field in our measurements varies between nominally identical flakes exfoliated from the same MnBi2Te4 crystal, whereas the magnitude of the reflection circular dichroism (RCD) at $H=0$ remains constant. This observation is illustrated in Fig. S1, in which we compare two flakes placed on the same substrate and measured in the same cooldown. While the zero-field RCD values are identical within experimental error, the coercive fields and the shapes of the hysteresis differ.
We attribute this difference in coercivity to variations in local surface conditions, such as strain introduced during exfoliation or subtle differences in surface oxidation. We know that the Hamiltonian parameters of the surface spin (exchange coupling, anisotropy) are different from those of the bulk, and as we point in the main text this difference is crucial for magnetic-field tuning of AFM domains. The difference between the Hamiltonian properties at the surface and the bulk can of course be influenced by details of the surface, leading to the different coercivity. However, surface oxidation in MnBi2Te4 is self-limiting and confined to the topmost atomic layer [32], and the optical penetration depth far exceeds this layer thickness. Consequently, while the surface environment can affect the switching field of the antiferromagnetic domains, it cannot account for the robust and reproducible RCD magnitude at $H=0$.
[…]
To further test whether surface modifications could mimic our results, we simulated RCD spectra for three scenarios: (a) pristine A-type AFM with domain reversal; (b) A-type AFM with domain reversal, with a non-magnetic surface layer; (c) AFM with a ferromagnetic surface layer that switches independently of the bulk. Only scenarios (a,b), in which the AFM domain switches, reproduce the experimental finding that the two zero-field RCD magnitudes are equal in magnitude and opposite in sign (Fig. ED5). In particular, switching only a surface ferromagnetic layer yields asymmetric zero-field spectra, inconsistent with the data (Fig. S2).
Figure S2: Effect of surface modifications on zero-field RCD. Simulated RCD spectra for three scenarios: (a) pristine A-type AFM with domain reversal; (b) A-type AFM with domain reversal and a non-magnetic surface layer; and (c) A-type AFM with a ferromagnetic surface layer that switches independently of the bulk. Each panel shows RCD as a function of photon energy for the two AFM domain configurations sketched in the corner. Only scenarios (a) and (b), in which the AFM domain switches, reproduce the experimental observation that the two zero-field RCD spectra are equal in magnitude and opposite in sign. Switching only a surface ferromagnetic layer yields asymmetric zero-field spectra, inconsistent with the experiment.
These findings establish that the reproducible zero-field RCD signal arises from reversal of bulk AFM domains. While the coercive field is sensitive to details of the surface, this does not influence the magnitude of the RCD at $H=0$, which is a robust signature of the underlying bulk antiferromagnetic order.